# Towards Hierarchical Multi-Step Reward Models for Enhanced Reasoning in Large Language Models

### v3

@@ Abstract
Recent studies show that Large Language Models (LLMs) achieve strong reasoning capabilities through supervised fine-tuning or reinforcement learning. However, a key approach, the Process Reward Model (PRM), suffers from reward hacking, making it unreliable in identifying the best intermediate step. Moreover, the cost of annotating reasoning processes for reward modeling is high, making large-scale collection of high-quality data challenging. To address the first problem, we propose a novel reward model approach, Hierarchical Reward Model (HRM), which evaluates both individual and consecutive reasoning steps from fine-grained and coarse-grained level. HRM excels at assessing multi-step reasoning coherence, particularly in cases where a flawed step is later corrected through self-reflection. Furthermore, to address the inefficiency of autonomous annotating PRM training data via Monte Carlo Tree Search (MCTS), we introduce a lightweight and effective data augmentation strategy called Hierarchical Node Compression (HNC) based on node merging (combining two consecutive reasoning steps into one step) in the tree structure. By applying HNC to MCTS-generated reasoning trajectories, we enhance the diversity and robustness of HRM training data while introducing controlled noise with minimal computational overhead. Empirical results on the PRM800K dataset demonstrate that HRM, in conjunction with HNC, achieves superior stability and reliability in evaluation compared to PRM. Furthermore, cross-domain evaluations on MATH500 and GSM8K dataset confirm HRM’s superior generalization and robustness across diverse reasoning tasks.
@@ 1 Introduction
As the scale of parameters in Large Language Models (LLMs) continues to grow Anil et al. (2023); Achiam et al. (2023); Grattafiori et al. (2024); Yang et al. (2024a), their general capabilities have significantly improved, surpassing human performance in various generative tasks such as text comprehension and data generation Wang et al. (2024a). However, the upper bound and inherent limitations of LLMs in reasoning-intensive tasks—such as mathematical and logical reasoning—remain an open question Cobbe et al. (2021); Lightman et al. (2023); Uesato et al. (2022); Wang et al. (2023); Luo et al. (2024); Amrith Setlur (2025); Wang et al. (2024b, c). Recent approaches, such as Chain-of-Thought (CoT)Wei et al. (2022) and Tree-of-Thought (ToT)Shunyu Yao (2023), have significantly enhanced reasoning performance. Despite these advancements, most CoT models lack mechanisms to detect and correct intermediate reasoning errors, resulting in continued propagation of mistakes throughout the reasoning process. Meanwhile, ToT methods do not inherently verify every intermediate step or guarantee retrieval of the optimal reasoning trajectory, which can limit its reliability in complex problem-solving scenarios.
To mitigate these limitations, recent efforts have focused on reward mechanisms that guide LLMs more effectively. There are two primary approaches to enhance the reasoning capabilities of LLMs from the perspective of "how to reward LLMs": the Outcome Reward Model (ORM) Lightman et al. (2023); Uesato et al. (2022); Guo et al. (2025); Shao et al. (2024) and the Process Reward Model (PRM) Lightman et al. (2023); Uesato et al. (2022). Each comes with its own limitations. ORM suffers from delayed feedback and credit assignment issues, making it difficult to pinpoint which reasoning steps contribute to the final answer Lightman et al. (2023); Uesato et al. (2022). PRM, in contrast, provides finer-grained supervision by evaluating reasoning step by step. However, most PRM methods are model-based and are prone to reward hacking Amodei et al. (2016); Di Langosco et al. (2022); Everitt et al. (2017), where models exploit reward signals rather than genuinely improving reasoning, undermining reliability in complex tasks. Moreover, the high annotation cost associated with PRM makes large-scale deployment challenging.
@@
Granularity Whole Process Single Step Few Consecutive Steps
@@
In this paper, we focus on addressing the limitations of PRM. To mitigate the impact of reward hacking in PRM, we propose the Hierarchical Reward Model (HRM). Traditional PRM penalizes an incorrect single step without considering potential corrections in subsequent reasoning. In contrast, HRM assesses reasoning coherence across multiple steps, allowing the reward model to identify and incorporate later steps that rectify earlier errors, leading to a more robust and reliable evaluation. Table 1 compares the difference between ORM, PRM and HRM. The PRM800K Lightman et al. (2023) dataset comprises manually annotated reasoning trajectories, which serve as the foundation for training ORM, PRM, and HRM. We subsequently assess the performance of Qwen2.5-72B-Math-Instruct Yang et al. (2024b) as the policy model by employing the Best-of-N Search strategy across ORM, PRM, and HRM. Experimental results (as shown in Table 3) demonstrate that HRM is the most robust to reward hacking. The policy model with HRM maintains stable performance, with accuracy stabilizing at 80% as $N$ increases. In contrast, policy models with PRM and ORM exhibit significant performance fluctuations, with accuracy degrading as $N$ grows.
@@
To fully exploit the capabilities of Monte Carlo Tree Search (MCTS), we introduce a data augmentation framework termed Hierarchical Node Compression (HNC), which consolidates two consecutive nodes from different depths into a single node. This approach effectively expands the training dataset while maintaining minimal computational overhead and enhancing label robustness through controlled noise injection. After evaluating HNC in the auto-annotation process by MCTS on the PRM800K dataset, we find that fine-tuned HRM achieves more robust scoring within PRM800K dataset and exhibits strong generalization across other domains, including GSM8K Cobbe et al. (2021) and MATH500 Lightman et al. (2023) dataset, outperforming PRM in robustness and consistency.
Our main contribution are as follows:
@@
  • We propose the Hierarchical Reward Model (HRM), which enhances multi-step reasoning coherence and self-reflection by evaluating both fine-grained reasoning steps and coarse-grained reasoning trajectories. We validate HRM’s robustness on the PRM800K dataset using manually annotated data.
  • We introduce Hierarchical Node Compression (HNC) in MCTS for autonomous reasoning process annotation, refining existing reasoning trajectories by generating additional step-wise labels with controlled noise to improve score robustness. Our method can greatly expand the HRM training dataset with minimal overhead.
@@
  • Additionally, we enhance the policy model through fine-tuning on high-quality reasoning trajectories filtered from MCTS, further improving its reasoning performance. We train HRM on PRM800K dataset with auto-annotated reasoning data from HNC and MCTS, demonstrating its improved robustness over PRM. Furthermore, HRM trained on PRM800K dataset exhibits superior reasoning consistency and generalization across GSM8K and MATH500 dataset, outperforming PRM.
@@ 2.1 RLHF
Reinforcement Learning with Human Feedback (RLHF) Ouyang et al. (2022) is a widely used framework for optimizing LLMs by incorporating human feedback signals. The core idea of RLHF is to use a Reward Model (RM) to distinguish between high-quality and low-quality responses and then optimize the LLM using Proximal Policy Optimization (PPO) Schulman et al. (2017). Ouyang et al. (2022) first apply RLHF to train InstructGPT Ouyang et al. (2022), aligning the model’s outputs with human preferences.
From the perspective of reward design, there are two main approaches: Outcome Reward Models (ORM) Cobbe et al. (2021); Lightman et al. (2023); Wang et al. (2023) and Process Reward Models (PRM) Lightman et al. (2023); Uesato et al. (2022); Luo et al. (2024); Zhang et al. (2024). ORM assigns rewards based on the whole output, while PRM evaluates intermediate reasoning steps to provide more fine-grained supervision. These reward mechanisms directly impact how LLMs learn to reason and optimize their outputs.
@@ 2.2 ORM
ORM suffers from delayed feedback and the credit assignment problem. Since rewards are only provided at the final outcome, ORM struggles to discern which intermediate steps contribute to success or failure Cobbe et al. (2021); Lightman et al. (2023); Wang et al. (2023). This delayed feedback limits learning efficiency, making it harder to optimize critical decision points. Additionally, ORM is prone to spurious reasoning Cobbe et al. (2021); Lightman et al. (2023); Wang et al. (2023), where the model arrives at the correct answer despite flawed intermediate steps, reinforcing suboptimal reasoning patterns. However, DeepSeek-R1Guo et al. (2025) integrates a rule-based ORM within the Group Relative Policy Optimization (GRPO) algorithmShao et al. (2024), demonstrating that rule-based reward mechanisms, rather than score-based reward models, can effectively guide LLMs toward generating long-chain-of-thought (long-CoT) reasoning and self-reflection, ultimately enhancing their reasoning abilities.
@@ 2.3 PRM
In the LLM domain,  Wang et al. (2024d) find that when an LLM is used as a verifier to assess the quality of multiple outputs, its ranking can be easily manipulated simply by changing the order of candidates in the context.  Wen et al. (2025) demonstrate that RLHF can make AI models more persuasive to human evaluators without necessarily improving their correctness, leading to higher human approval of incorrect answers and increased evaluation error rates.  Pan et al. (2024) analyze In-Context Reward Hacking (ICRH), where an AI system engaged in a feedback loop optimizes a proxy objective—such as a writing assistant maximizing “user satisfaction scores”—and inadvertently learns to generate exaggerated claims or misinformation if these result in higher perceived satisfaction.
Furthermore, the annotation process required for training PRM models is prohibitively expensive Lightman et al. (2023); Uesato et al. (2022), making large-scale implementation impractical. To address this issue, Monte Carlo Tree Search (MCTS) has been proposed as an autonomous method for annotating reasoning trajectories Luo et al. (2024); Zhang et al. (2024). While MCTS reduces the need for human annotation, it incurs substantial computational costs due to the extensive simulations required for the MC-Score to achieve relative convergence. Moreover, in MCTS, the computational cost increases significantly as the depth and breadth of the search tree expand. To mitigate this, constraints are imposed on both tree height and width, limiting the number of simulation steps and thereby reducing the diversity of generated reasoning data.
@@ 3.1 Hierarchical Reward Model
PRM Lightman et al. (2023); Uesato et al. (2022) focuses on fine-grained, step-wise reasoning, while ORM Lightman et al. (2023); Uesato et al. (2022) evaluates reasoning as a whole. To integrate the strengths of both, we propose the Hierarchical Reward Model (HRM), which evaluates individual reasoning steps while also ensuring multi-step coherence by analyzing consecutive steps. The training dataset for HRM consists of consecutive reasoning sequences spanning from step 1 to $N$, as illustrated in Fig. 1 and Section 4.1. HRM training data is a superset of PRM training data, constructed by merging consecutive reasoning steps. Specifically, while PRM training data only contains individual steps with their respective reward scores, HRM training data additionally incorporates merged step pairs to enhance reasoning coherence and improve self-reflection in error correction. Formally, let $D$ represent the training dataset, $N$ denote the total number of reasoning steps in a sequence, $s_{i}$ be the $i$-th reasoning step, and $R(\cdot)$ be the reward function that assigns a score to a step. The training datasets for PRM and HRM are defined as:
$D_{\text{PRM}}=\{(s_{i},R(s_{i}))\mid 1\leq i\leq N\},$
$D_{\text{HRM}}=D_{\text{PRM}}\cup\{(s_{i}+s_{i+1},R(s_{i}+s_{i+1}))\mid 1\leq i<N\}.$
HRM is designed with two primary objectives: (1) capturing both fine-grained and coarse-grained reasoning consistency, and (2) enabling self-reflection and error correction. Unlike PRM, which terminates evaluation upon encountering an error, HRM assesses whether subsequent steps rectify earlier mistakes, treating them as a cohesive unit rather than isolated errors.
@@
Figure 2: Illustration of the MCTS-based automated reasoning annotation process. The left side depicts a tree structure where each node represents a reasoning step, simulated using the Tree-of-Thought approach with MCTS. The right side visualizes the assigned scores for each step in the reasoning tree.
@@ 3.2 Hierarchical Node Compression in MCTS
Although process supervision enhances the reasoning capabilities of policy models, the cost of human-annotated supervision is prohibitively high Lightman et al. (2023); Uesato et al. (2022). To address this, autonomous annotation methods based on MCTS have been proposed Luo et al. (2024); Zhang et al. (2024). Fig. 2 illustrates the process of automatic reasoning annotation using MCTS. Given a ground truth and a corresponding question, MCTS generates multiple possible reasoning paths by simulating different step-by-step solutions. Each node in the search tree represents a reasoning step, and its score is calculated based on the proportion of correct trajectories in its subtree, reflecting the likelihood that the reasoning path is valid. However, these methods demand substantial computational resources, as achieving reliable estimates of intermediate reasoning step scores requires a sufficiently deep and wide search tree to reduce variance and mitigate bias; otherwise, the estimates may remain unreliable. This exponential growth in complexity makes large-scale implementation challenging.
To address the limitations of autonomous process annotation, we propose a data augmentation method called Hierarchical Node Compression (HNC), which maximally leverages MCTS-generated data. The key idea is to merge two consecutive nodes, each corresponding to a reasoning step, into a single node, thereby creating a new branch with minimal computational overhead.
As illustrated in Fig. 3, HNC assumes that each node has a sufficiently large number of child nodes. By randomly removing or merging consecutive nodes, it introduces controlled noise, enhancing the robustness of MCTS-based scoring. Before HNC, each child node contributes $\frac{1}{N}$ to the total score. After HNC removes a random node, the remaining child nodes redistribute their weights to $\frac{1}{N-1}$, increasing their individual influence. Since child nodes are independent and identically distributed (i.i.d.) from the parent’s perspective, the expectation of the parent score remains unchanged. However, the variance increases from $\frac{\sigma^{2}}{N}$ to $\frac{\sigma^{2}}{N-1}$, introducing controlled noise that enables data augmentation at an extremely low computational cost. When $N$ is sufficiently large, this variance change remains moderate while still facilitating effective data augmentation.
Figure 3: Illustration of Hierarchical Node Compression. The left part represents the original MCTS data annotation structure, while the right part shows the transformed MCTS structure after applying HNC.

### v4

@@ Abstract
Large Language Models (LLMs) have demonstrated strong mathematical reasoning abilities through supervised fine-tuning and reinforcement learning. However, existing Process Reward Models (PRMs) are vulnerable to reward hacking and require expensive, large-scale annotation of reasoning steps, limiting their reliability and scalability. To address the first problem, we propose a novel reward model approach, Hierarchical Reward Model (HRM), which evaluates both individual and consecutive reasoning steps from fine-grained and coarse-grained level. HRM excels at assessing multi-step mathematical reasoning coherence, particularly in cases where a flawed step is later corrected through self-reflection. Furthermore, to address the inefficiency of autonomously annotating PRM training data via Monte Carlo Tree Search (MCTS), we propose a lightweight data augmentation strategy, Hierarchical Node Compression (HNC), which merges consecutive reasoning steps within the tree structure. Applying HNC to MCTS-generated reasoning trajectories increases the diversity and robustness of HRM training data, while introducing controlled noise with minimal computational overhead. Empirical results on the PRM800K dataset demonstrate that HRM, in conjunction with HNC, achieves superior stability and reliability in evaluation compared to PRM. Furthermore, cross-domain evaluations on MATH500 and GSM8K dataset confirm HRM’s superior generalization and robustness across diverse mathematical reasoning tasks.
Towards Hierarchical Multi-Step Reward Models for Enhanced Reasoning in Large Language Models
Teng Wang1, Zhangyi Jiang2, Zhenqi He1, Hailei Gong5thanks: Corresponding author, Shenyang Tong1, Wenhan Yang1, Zeyu Li4, Yanan Zheng3, Zifan He2, Zewen Ye6, Shengjie Ma7 Jianping Zhang8 1 the University of Hong Kong, 2 Peking University, 3 National University of Singapore, 4 Georgia Institute of Technology, 5 Tsinghua University, 6 Zhejiang University, 7 Renmin University of China, 8 Meta wt0318@connect.hku.hk
@@ 1 Introduction
Granularity (Training Data) Whole Process Single Step Few Consecutive Steps
@@
As the scale of parameters in LLMs continues to grow Anil et al. (2023); Achiam et al. (2023); Grattafiori et al. (2024); Yang et al. (2024a), their general capabilities have significantly improved, surpassing human performance in various generative tasks such as text comprehension and data generation Wang et al. (2024a). However, the upper bound and inherent limitations of LLMs in reasoning-intensive tasks—such as mathematical reasoning—remain an open question Cobbe et al. (2021); Lightman et al. (2023); Uesato et al. (2022); Wang et al. (2023); Luo et al. (2024); Amrith Setlur (2025); Wang et al. (2024c, b). Recent approaches, such as Chain-of-Thought (CoT) Wei et al. (2022) and Tree-of-Thought (ToT) Shunyu Yao (2023), have significantly enhanced reasoning performance. Despite these advancements, most CoT models lack mechanisms to detect and correct intermediate reasoning errors, resulting in continued propagation of mistakes throughout the reasoning process. Meanwhile, ToT methods do not inherently verify every intermediate step or guarantee retrieval of the optimal reasoning trajectory, which can limit its reliability in complex problem-solving scenarios.
To mitigate these limitations, recent efforts have focused on reward mechanisms that guide LLMs effectively. There are two primary approaches to enhance the reasoning capabilities of LLMs from the perspective of "how to reward LLMs": the Outcome Reward Model (ORM) Lightman et al. (2023); Uesato et al. (2022); Guo et al. (2025); Shao et al. (2024) and the Process Reward Model (PRM) Lightman et al. (2023); Uesato et al. (2022). Each comes with its own limitations. ORM suffers from delayed feedback and credit assignment issues, making it difficult to pinpoint which reasoning steps contribute to the final answer Lightman et al. (2023); Uesato et al. (2022). PRM, in contrast, provides finer-grained supervision by evaluating reasoning step by step. However, most PRM methods are model-based and are prone to reward hacking Weng (2024), where models exploit reward signals rather than genuinely improving reasoning, undermining reliability in complex tasks. Moreover, the high annotation cost associated with PRM makes large-scale deployment challenging.
@@
In this paper, we focus on addressing the limitations of PRM. To mitigate the impact of reward hacking in PRM, we propose the Hierarchical Reward Model (HRM). The term hierarchical highlights that, during training, HRM incorporates a hierarchical supervision signal by evaluating reasoning processes at both fine-grained (single-step) and coarse-grained (consecutive multi-step) levels. This layered approach enables HRM to capture both local and global coherence in reasoning. However, during inference, HRM remains step-wise: it assigns rewards to each reasoning step individually, the same as PRM. Traditional PRM penalizes an incorrect single step without considering potential corrections in subsequent reasoning. In contrast, HRM assesses reasoning coherence across multiple steps, allowing the reward model to identify and incorporate later steps that rectify earlier errors, leading to a more robust and reliable evaluation. Table 1 compares the difference between ORM, PRM and HRM.
While HRM can be applied to other structured reasoning domains, we focus on mathematical reasoning, the only domain with large-scale human-annotated process-level data for supervised reward modeling. The PRM800K Lightman et al. (2023) dataset comprises manually annotated reasoning trajectories, which serve as the foundation for training ORM, PRM, and HRM. We subsequently assess the performance of Qwen2.5-72B-Math-Instruct Yang et al. (2024b) as the policy model by employing the Best-of-N Search strategy across ORM, PRM, and HRM. Experimental results (as shown in Table 2) demonstrate that HRM yields the most stable and reliable reward evaluations. The policy model with HRM maintains stable performance, with accuracy stabilizing at 80% as $N$ increases. In contrast, policy models with PRM and ORM exhibit significant performance fluctuations, with accuracy degrading as $N$ grows.
@@
To fully exploit the capabilities of Monte Carlo Tree Search (MCTS) for automatic process annotation, we introduce a data augmentation framework termed Hierarchical Node Compression (HNC), which consolidates two consecutive nodes from different depths into a single node. This approach effectively expands the training dataset while maintaining minimal computational overhead and enhancing label robustness through controlled noise injection. After evaluating HNC in the auto-annotation process by MCTS on the PRM800K dataset, we find that fine-tuned HRM achieves more robust scoring within PRM800K dataset and exhibits strong generalization across other domains, including GSM8K Cobbe et al. (2021) and MATH500 Lightman et al. (2023) dataset, outperforming PRM in robustness and consistency.
Our main contributions are as follows:
@@
  • We propose the HRM, which leverages hierarchical supervision from training data at both single-step and multi-step levels, promoting coherence and self-correction in multi-step reasoning. We validate HRM’s robustness on the PRM800K dataset using manually annotated data.
  • We introduce HNC, a lightweight data augmentation approach for MCTS that substantially increases the diversity and robustness of HRM training data with minimal computational cost. Experiments show that HRM trained on the PRM800K dataset with auto-annotated data from HNC and MCTS demonstrates improved robustness over PRM. Furthermore, HRM exhibits superior reasoning consistency and generalization across GSM8K and MATH500, consistently outperforming PRM.
@@
  • Additionally, we enhance the policy model through fine-tuning on high-quality reasoning trajectories filtered from MCTS, further improving its reasoning performance.
@@ 2.1 RLHF
Reinforcement Learning with Human Feedback (RLHF) Ouyang et al. (2022) is a widely used framework for optimizing LLMs by incorporating human feedback. The core idea of RLHF is to use a Reward Model (RM) to distinguish between high-quality and low-quality responses and optimize the LLM using PPO Schulman et al. (2017).
From the perspective of reward design, there are two main approaches: ORM Cobbe et al. (2021); Lightman et al. (2023); Wang et al. (2023) and PRM Lightman et al. (2023); Uesato et al. (2022); Luo et al. (2024); Zhang et al. (2024). ORM assigns rewards based on the whole output, while PRM evaluates intermediate reasoning steps to provide more fine-grained supervision.
@@ 2.2 ORM
ORM suffers from delayed feedback and the credit assignment problem. Since rewards are only provided at the final outcome, ORM struggles to discern which intermediate steps contribute to success or failure Lightman et al. (2023). This delayed feedback limits learning efficiency, making it harder to optimize critical decision points. Additionally, ORM is prone to spurious reasoning Cobbe et al. (2021); Wang et al. (2023), where the model arrives at the correct answer despite flawed intermediate steps, reinforcing suboptimal reasoning patterns. However, DeepSeek-R1Guo et al. (2025) integrates a rule-based ORM within GRPOShao et al. (2024), demonstrating that rule-based reward, rather than score-based reward models, can effectively guide LLMs toward generating long-CoT reasoning and self-reflection, ultimately enhancing their reasoning abilities.
Figure 2: Illustration of the MCTS-based automated reasoning annotation process. The left side depicts a tree structure where each node represents a reasoning step, simulated using the ToT approach with MCTS. The right side visualizes the assigned scores for each step in the reasoning tree.
@@ 2.3 PRM
Furthermore, the annotation process required for training PRM is prohibitively expensive Lightman et al. (2023); Uesato et al. (2022), making large-scale implementation impractical. To address this issue, MCTS has been proposed as an autonomous method for annotating reasoning trajectories Luo et al. (2024); Zhang et al. (2024). Moreover, in MCTS, the computational cost increases significantly as the depth and breadth of the search tree expand. To mitigate this, constraints are imposed on both tree height and width, limiting the number of simulation steps and thereby reducing the diversity of generated reasoning data.
@@ 3.1 Hierarchical Reward Model
PRM provides fine-grained, step-wise supervision, whereas ORM evaluates reasoning holistically. To combine the advantages of both, we propose the Hierarchical Reward Model (HRM), which introduces hierarchical supervision by training on both single-step and consecutive multi-step reasoning sequences. HRM serves as an evaluation-oriented reward model that estimates the relative quality of reasoning trajectories, rather than directly optimizing the policy model itself. This hierarchical design enables HRM to capture both local accuracy and global coherence, yielding more stable and reliable reward evaluation in multi-step reasoning. The HRM training data extends PRM supervision by merging consecutive reasoning steps into hierarchical segments (from step 1 to $N$), as illustrated in Fig. 1, forming a superset of the PRM dataset.
Figure 3: Illustration of HNC. The left part represents the original MCTS data annotation structure, while the right part shows the transformed MCTS structure after applying HNC.
Formally, let $D$ represent the training dataset, $N$ denote the total number of reasoning steps in a sequence, $s_{i}$ be the $i$-th reasoning step, and $R(\cdot)$ be the reward function that assigns a score to a step. The training datasets for PRM and HRM are defined as:
$D_{\text{PRM}}=\left\{(s_{i},\,R(s_{i}))\mid 1\leq i\leq N\right\},$ (1)
$D_{\text{HRM}}=D_{\text{PRM}}$
$\quad\cup\left\{(s_{i}+s_{i+1},\,R(s_{i}+s_{i+1}))\mid 1\leq i<N\right\}.$ (2)
Hierarchical supervision aims to connect local correctness with global coherence by jointly evaluating individual steps and their interactions. In practice, HRM achieves this through two objectives: (1) capturing both fine-grained and coarse-grained consistency, and (2) enabling self-reflection and error correction. Unlike PRM, which evaluates each step independently and penalizes early mistakes, HRM considers whether later steps can repair earlier errors, treating them as a coherent reasoning process rather than isolated faults.
@@
This design preserves fine-grained step-level evaluation while mitigating the subjectivity of step segmentation: by merging neighboring steps during training, HRM smooths annotation noise from ambiguous boundaries and enhances robustness across reasoning granularities. Conceptually, this hierarchical supervision extends PRM in a principled way, integrating both local accuracy and global coherence. In this work, we instantiate HRM with two-step merging as the minimal yet effective form of hierarchical supervision, enabling controlled analysis while maintaining computational efficiency.
@@ 3.2 Hierarchical Node Compression in MCTS
Due to the prohibitively high cost of human-annotated supervision, autonomous annotation methods based on MCTS have been proposed. Fig. 2 illustrates the process of automatic reasoning annotation using MCTS. Given a ground truth and a corresponding question, MCTS generates multiple possible reasoning paths by simulating different step-by-step solutions. Each node in the search tree represents a reasoning step, and its score is calculated based on the proportion of correct trajectories in its subtree, reflecting the likelihood that the reasoning path is valid. However, these methods demand substantial computational resources, as achieving reliable estimates of intermediate reasoning step scores requires a sufficiently deep and wide search tree to reduce variance and mitigate bias; otherwise, the estimates may remain unreliable. This exponential growth in complexity makes large-scale implementation challenging.
To better leverage autonomous process annotation, we propose a data augmentation method called Hierarchical Node Compression (HNC). The key idea is to merge two consecutive nodes, each corresponding to a reasoning step, into a single node, thereby creating a new branch with minimal computational overhead.
As shown in Fig. 3, HNC assumes that each node has a sufficiently large number of child nodes. By randomly merging consecutive nodes, it introduces controlled noise, enhancing the robustness of MCTS-based scoring. Before HNC, each child node contributes $\frac{1}{N}$ to the total score. After HNC removes a random node, the remaining child nodes redistribute their weights to $\frac{1}{N-1}$, increasing their individual influence. Since child nodes are independent and identically distributed from the parent’s perspective, the expectation of the parent score remains unchanged. However, the variance increases from $\frac{\sigma^{2}}{N}$ to $\frac{\sigma^{2}}{N-1}$, introducing controlled noise that enables data augmentation at an extremely low computational cost. When $N$ is sufficiently large, this variance change remains moderate while still facilitating effective data augmentation.